# EEG-DLite: Dataset Distillation for Efficient Large EEG Model Training

### v1

@@ Introduction
Brain-computer interfaces (BCIs) enable direct communication between the human brain and external devices, unlocking a wide range of applications in neurorehabilitation (mtp-bci), assistive technologies (Tang2024), and cognitive monitoring (zhou2025brainfoundationmodelssurvey). Among the various modalities used in BCIs, electroencephalography (EEG) is the most widely adopted due to its non-invasiveness and high temporal resolution. However, EEG signals are inherently noisy, high-dimensional, and subject to significant inter-subject variability, which presents challenges for learning robust and generalizable representations (WangNth2025; carzaniga2025the).
To address these challenges, recent advances have introduced EEG foundation models, which are large-scale neural networks pre-trained on extensive unlabeled EEG datasets using self-supervised objectives (zhou2025brainfoundationmodelssurvey). These models, often built on convolutional backbones or Transformer-based architectures, can exceed 400 million parameters. Typically, they are first pre-trained on large-scale EEG data from diverse BCI tasks and later fine-tuned for specific downstream applications. These models demonstrate competitive performance across various downstream tasks while maintaining a unified architectural framework (jiang2024largebrainmodellearning; eegpt; wang2025cbramod). While effective, this pre-training paradigm is computationally expensive, often requiring significant GPU time, storage, and energy (jiang2024largebrainmodellearning). The high cost also limits the feasibility of further exploration into parameter optimization and neural architecture search (yu2023datasetdistillationcomprehensivereview).
@@
Recent advances in computer vision have shown that it is possible to significantly reduce the computational burden of large-scale model training by distilling extensive datasets into compact, information-rich subsets (zhang2024m3ddatasetcondensationminimizing; joshi2024datasetdistillationknowledgedistillation; killamsetty2021gradmatchgradientmatchingbased). Motivated by the broader goal of data efficient pre-training, we turn our focus to EEG data, which presents unique and domain-specific challenges that demand tailored solutions. First, EEG signals exhibit a low signal-to-noise ratio (SNR), as neural activity is easily obscured by artifacts from eye movements, muscle activity, and external interference (9441341). Even after preprocessing, residual noise can persist and adversely impact model learning. Second, EEG recordings often contain significant redundancy; temporally adjacent segments may capture overlapping or repetitive patterns, offering limited new information (amin2016novel). These characteristics make EEG fundamentally different from vision data in structure and signal quality, necessitating data-efficient strategies that directly address its inherent noise and redundancy. Despite these challenges, little is known about how the volume and composition of pre-training EEG data influence the generalization ability of foundation models, leaving a critical gap in the development of scalable and effective large-scale EEG modeling.
To overcome these limitations, we propose EEG-D2, a data distillation framework for data efficient pre-training of large EEG foundation models. EEG-D2 systematically prunes large, unlabeled EEG datasets that are collected using varied channel montages into smaller, more representative subsets. Due to the high-dimensionality and low signal-to-noise ratio of EEG signals, instead of selecting samples directly using EEG signals, our approach begins by encoding EEG segments into compact latent representations using a self-supervised autoencoder. Based on these representations, we apply two key techniques:
@@
These steps produce a significantly smaller dataset that still captures the diversity and structure necessary for effective model training. In our experiments, training on only 5% of a 2,500-hour EEG dataset distilled using EEG-D2 achieves comparable, or even superior, performance on several downstream tasks compared to training on the full dataset, as shown in Figure 1. Additionally, this reduces GPU pre-training time from 30 hours to just 2 hours, under the same hardware conditions.
To our knowledge, this is the first study to explore data distillation for physiological signals like EEG in the context of large foundation models. We also perform systematic comparisons of generative and selection-based distillation approaches, providing new insights into what makes pre-training data effective in EEG modeling.
@@
  • We present the first comparison between synthetic data generation and selection-based distillation on EEG. The results show the challenges EEG pose on the generation approach and highlights the effectiveness of a core-set selection approach.
@@
4: Compute outlier scores: $\prod_{k=1}^{d}\log(1/(p_{k}(z_{ik})+\alpha))$
@@
$\min_{\bm{\mu}\subset\mathcal{Z}^{\prime}}\max_{z\in\mathcal{Z}^{\prime}}\min_{\mu_{k}\in\bm{\mu}}\|z-\mu_{k}\|_{2}^{2}$
7: Return subset $\mathcal{S}\subset\mathcal{X}$
@@
Figure 2: Overview of the proposed framework. In the first stage, an auto-encoder is trained with self-supervised learning to project each EEG segment $X$ into a lower-dimensional latent representation $\mathbf{z}$. In the second stage, out-of-distribution scores are computed for all samples, and the top $\tau\%$ with the highest scores are excluded. Finally, $\eta\%$ of the datasets are selected via diversity sampling to retain only the most informative data.
@@
EEG foundation models have been an emerging paradigm for decoding and modeling brain activity using deep neural networks pretrained on large-scale EEG recordings (zhou2025brainfoundationmodelssurvey; lai2025simplerevieweegfoundation). Typical architectures adopt self-supervised learning (SSL) objectives and employ CNNs, transformers, or hybrid architectures to extract generalizable neural representations. These models have shown strong performance on downstream tasks such as cognitive state decoding and clinical classification (jiang2024largebrainmodellearning; eegpt). However, training them requires massive datasets, often exceeding millions of EEG segments, which poses significant computational and storage burdens. More critically, existing efforts largely overlook the influence of training data composition and selection on model performance. While prior work has focused on architectural innovation and transfer learning strategies, few have explored how data quality, representativeness, and redundancy affect the foundation model’s generalization ability. Our work addresses this gap by introducing a data-centric framework that selects diverse, informative EEG segments to reduce training costs without compromising generalization. By emphasizing principled data selection, we aim to enhance both the scalability and explainability of large EEG models, bridging the methodological gap between neuroscience and AI through efficient and standardized data curation.
@@
Data distillation refers to the process of constructing a compact and informative subset of training data that allows a model to achieve comparable performance to that trained with the full dataset, while significantly reducing computational costs. Existing distillation methods generally fall into two main categories: synthetic data generation and core-set selection (joshi2024datasetdistillationknowledgedistillation; sener2018active). Synthetic data generation focuses on generate samples that simulate the learning behavior or statistical properties of the original dataset. They have gained popularity in many computer vision tasks due to the flexibility and expressive capacity (ding2024time). These methods often use optimization objectives such as performance matching, parameter matching, or distribution matching (yu2023datasetdistillationcomprehensivereview; zhang2024m3ddatasetcondensationminimizing). However, they typically rely on bi-level optimization, which involves nested optimization loops, making them computationally intensive. Core-set selection, in contrast, selects real and representative subset from the training data (sener2018active). They are more stable and interpretable since they work with real data, but their selection process becomes increasingly expensive as both data size and dimensionality grow. This is particularly problematic for EEG signals, which are high-dimensional, noisy, and vary across subjects. To address the computational burden, we adopt SSL to train a lightweight autoencoder that compresses EEG segments into low-dimensional latent representations, enabling efficient selection (balestriero2023cookbookselfsupervisedlearning).
@@
This section elaborates the details of the proposed framework, EEG-DLite, which is a model-agnostic EEG data distillation framework. It is designed to select a compact yet representative subset from massive unlabeled data for efficient pretraining. As illustrated in Figure 2 and Algorithm 1, the framework consists of three main steps: (1) a self-supervised multi-view autoencoder compresses high-dimensional EEG signals into low-dimensional latent representations; (2) an outlier detection module filters out anomalous or low-quality samples that may degrade diversity; and (3) a $k$-center selection algorithm identifies a diverse set of EEG segments that best represent the original dataset.
@@
To handle the high dimensionality of EEG signals, we train a lightweight autoencoder with SSL that maps raw EEG segments into compact latent representations without requiring labeled data (VAEEG; zhang2023selfsupervisedtimeseriesrepresentation). This architecture is intentionally decoupled from any specific foundation model to ensure flexibility and broad compatibility.
@@
The SSL objective includes two components: a reconstruction loss $\mathcal{L}_{\text{Rec}}$ and an inter-instance discrimination loss $\mathcal{L}_{\text{IDC}}$. The reconstruction loss minimizes the mean squared error between input patches $\bm{x}_{i}$ with length $L$ and reconstructions $\bm{x}_{i}^{\prime}$, ensuring the encoder can accurately encode of neural signal content. The loss of each sample $X$ is defined as
@@
$\mathcal{L}_{\text{IDC}}$ penalizes excessive token similarity between samples in one batch, encouraging feature diversity (chen2021iceinterinstancecontrastiveencoding). Specifically, denoting $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$} the $i$-th token before and after the encoder, the network projects them into the same embedding space with two separate projectors $g_{1}$ and $g_{2}$ and penalizes the cosine similarity between an original token and the encoded tokens in other samples, as described below,
@@
Table 1: Performance comparison across different distillation ratios ($\eta$) on SEED-V and MoBI datasets. $\bm{\kappa}$ refers to Cohen’s Kappa coefficient. Bold values indicate the best performance within certain distillation ratio, excluding the full data case. DS refers to the diversity sampling. The symbol indicates the result is significantly better ($p<0.05$) than the corresponding Random baseline at the same $\eta$ based on Mann–Whitney U test.
@@
PCA + DS 1 $22.59\pm 0.37$ $18\pm 0.50 22.$$2.16\pm 0.41$ $356\pm 0.004 0.0$$046\pm 0.005$ $14.90\pm 0.03$
5 $31.01\pm 0.40$ $13.35\pm 0.48$ $31.24\pm 0.37$ $0.534\pm 0.004$ $0.262\pm 0.000$ $13.32\pm 0.06$
10 $36.26\pm 0.31$ $20.09\pm 0.39$ $36.52\pm 0.29$ $\bm{0.541\pm 0.002}$ $0.265\pm 0.003$ $13.26\pm 0.02$
25 $38.61\pm 0.22$ $23.45\pm 0.28$ $39.00\pm 0.22$ $0.546\pm 0.001^{\dagger}$ $0.276\pm 0.004^{\dagger}$ $13.15\pm 0.03^{\dagger}$
Proposed 1 $\bm{29.65\pm 0.26}^{\dagger}$ $\bm{11.34\pm 0.32}^{\dagger}$ $\bm{29.66\pm 0.28}^{\dagger}$ $\bm{0.506\pm 0.007}^{\dagger}$ $\bm{0.225\pm 0.010}^{\dagger}$ $\bm{13.61\pm 0.10}^{\dagger}$
5 $\bm{38.59\pm 0.15}^{\dagger}$ $\bm{23.08\pm 0.19}^{\dagger}$ $\bm{38.90\pm 0.16}^{\dagger}$ $\bm{0.550\pm 0.001}^{\dagger}$ $\bm{0.283\pm 0.001}^{\dagger}$ $\bm{13.15\pm 0.07}^{\dagger}$
10 $\bm{39.07\pm 0.14}^{\dagger}$ $\bm{23.90\pm 0.19}^{\dagger}$ $\bm{39.47\pm 0.15}^{\dagger}$ $\bm{0.541\pm 0.002}$ $\bm{0.277\pm 0.014}$ $\bm{13.07\pm 0.04}^{\dagger}$
25 $\bm{39.65\pm 0.19}^{\dagger}$ $\bm{24.72\pm 0.22}^{\dagger}$ $\bm{40.07\pm 0.16}^{\dagger}$ $\bm{0.550\pm 0.003}^{\dagger}$ $\bm{0.281\pm 0.005}^{\dagger}$ $\bm{13.14\pm 0.06}^{\dagger}$
Full data 100 $40.95\pm 0.62$ $26.13\pm 0.75$ $41.20\pm 0.57$ $0.538\pm 0.010$ $0.288\pm 0.003$ $12.25\pm 0.03$
@@
Each sample $X$ can be assigned with an out-of-distribution (OOD) score, defined as Equation 3. The top $\tau$ percent samples are excluded from the subsequent diverse selection step to prevent degrading the diversity and quality of the distilled subset.
$\text{OOD}(X)=\prod_{i=1}^{d}\log\frac{1}{p_{i}(x_{i})+\alpha}$ (3)
@@
Table 2: Performance comparison across distillation ratios ($\eta$) on TUEV and TUAB datasets. $\bm{\kappa}$ refers to Cohen’s Kappa coefficient. Bold values indicate the best performance within certain distillation ratio, excluding the full data case. DS refers to the diversity sampling. The symbol indicates the result is significantly better ($p<0.05$) than the corresponding Random baseline at the same $\eta$ based on Mann–Whitney U test.
@@
The proposed framework utilizes the selection-based distillation method. We formalize the process of selecting the most representative EEG samples as the coreset construction problem. Formally, let $\mathcal{Z}\subset\mathbb{R}^{N\times d}$ be the EEG compressed pre-training datasets, where $N$ is the sample size, $d$ is the dimension of compressed representations. The objective is to select $k$ points from $\mathcal{Z}$ such that every data point $\mathbf{z}$ is closest to one selected center, and the largest distance between any point to its closest center is minimized, which is described as
@@
This problem is NP-hard, so we utilize greedy approximation solver proposed in (sener2018active) to iteratively select points that maximize the minimum distance to previously chosen k-centers, of which the time complexity is $\mathcal{O}(k\times N\times d)$.
Figure 3: UMAP visualization of compressed EEG segments after OOD removal and diversity sampling on SEED-sleep (seed-sleep).
@@
This section elaborates the baselines, implementation details, the setup during pre-training and fine-tuning stages, followed by the results and in-depth analysis.
@@
PCA (Anuragi2024; Fujiwara2020) reduces EEG dimensionality by projecting data onto top principal components. We use Incremental PCA to scale efficiently and assess how it compares with SSL-based embeddings in representing EEG diversity.
M3D (zhang2024m3ddatasetcondensationminimizing) generates synthetic data by minimizing embedding discrepancies using random, untrained networks. Although it is lightweight and computationally feasible at low distillation ratios, it still becomes impractical at 25 percent distillation ratio due to excessive GPU memory demands.
@@
The input signals from both the potential and spectral domains are independently normalized and segmented into 20 non-overlapping patches. Then, the neural compressor is trained for 50 epochs using the Adam optimizer, with the learning rate of 0.001 and gradient clipping applied at a maximum norm of 5.0. A scheduler is used to reduce the learning rate every 10 epochs with a decay factor of 0.5. The training objective, described in Equation 2, incorporates an instance discrimination loss term $\mathcal{L}_{\text{IDC}}$, weighted by $\beta=0.0001$. The encoder consists of 8 self-attention layers with 8 heads per layer, while the decoder contains 2 transformer layers, each with 8 attention-heads as well. Each EEG segment $X$ is projected into a 64-dimensional embedding space.
@@
We use the same pre-training datasets as in LaBraM (jiang2024largebrainmodellearning), which includes 32 datasets over 2,500 hours of EEG data collected from multiple tasks (e.g., emotion recognition, motor imagery, etc.). These datasets have diverse channel montages. Each EEG sample is a segment into $X\in\mathbb{R}^{c\times T}$, where $c$ is the number of channels and $T$ is the temporal length (either 4 or 8 seconds), with a fixed stride of 4 seconds. All data are band-pass filtered within the range [0.1, 75] Hz and resampled to 200Hz. During the distillation process, each dataset is distilled independently.
@@
To assess the effectiveness of each distillation method, we evaluate the pretrained foundation models on diverse types of downstream tasks. To perform rigorous comparison, each method distills the original pretraining dataset at multiple compression ratios ($\eta=1\%,5\%,10\%,25\%$) first, and the resulting distilled subsets are used for pretraining. Finally, the pretrained models are fine-tuned separately on four representative downstream tasks that cover both classification and regression settings. We follow the identical evaluation protocol established in LaBraM to ensure direct comparability. All experiments use consistent hyperparameters during pretraining and fine-tuning. Results are averaged across five random seeds, and standard deviations and results’ significance are reported in Table 1 and 2. Below are the details of downstream tasks:
  • TUEV (10.3389/fnins.2016.00196): Six-class classification of EEG events.
  • TUAB (10.3389/fnins.2016.00196): Binary classification of normal versus abnormal EEG signals.
  • SEED-V (SEED-V): Five-class emotion recognition based on EEG.
  • MoBI (mobi): Continuous regression of bilateral lower-limb joint angles from EEG during walking.
@@
The foundation model retains comparable performance with  5% of the pretraining data.
As the dataset percentage increases, the performance on all the downstream tasks demonstrate improvements, indicating enhanced model generalization with more training data, and the performance change is also aligned with the experiments reported in (jiang2024largebrainmodellearning). Tables 1 and 2 indicate a consistent pattern of data redundancy in large-scale EEG pretraining datasets. Specifically, we observe that using only  5% of the pre-training data selected by EEG-DLite achieves performance close to the upper bound. In contrast, the random selection baseline requires around 25% of the data to reach a similar level. These results suggest that a substantial portion of the data has limited impact on model performance, highlighting the potential for more efficient training and deployment through targeted data selection.
@@
Method $\bm{\eta}$ (%) Acc. F1 $\bm{\kappa}$
Random 50 53.36 53.99 28.94
25 52.79 52.06 27.95
PCA + DS 50 51.76 48.60 27.37
25 51.56 49.92 26.22
Proposed ($\tau=0$) 50 54.07 52.82 29.90
25 54.56 55.11 29.11
Proposed ($\tau=1\%$) 50 56.56 56.71 33.19
25 55.25 55.68 31.28
Full Data 100 54.56 55.36 30.79
@@
To further evaluate the effectiveness of EEG-DLite, we conduct a pilot study on SEED dataset (zheng2015investigating; duan2013differential) using small network like EEGNet in the cross-subject supervised learning setting. The results in Table 3 show that EEG-DLite achieves the highest performance when trained on only 25% of the data, even outperforming models trained on the full dataset. We also observe the model is sensitive to the presence of OOD samples in supervised setting.

### v2

@@ Introduction
Brain-computer interfaces (BCIs) enable direct communication between the human brain and external devices, unlocking a wide range of applications in neurorehabilitation (Wang et al. 2023), assistive technologies (Tang et al. 2024), and cognitive monitoring (Zhou et al. 2025). Among the various modalities used in BCIs, electroencephalography (EEG) is the most widely adopted due to its non-invasiveness and high temporal resolution. However, EEG signals are inherently noisy, high-dimensional, and subject to significant inter-subject variability, which presents challenges for learning robust and generalizable representations (Wang-Nöth et al. 2025; Carzaniga et al. 2025).
To address these challenges, recent advances have introduced EEG foundation models, which are large-scale neural networks pre-trained on extensive unlabeled EEG datasets using self-supervised objectives (Zhou et al. 2025). These models, often built on convolutional backbones or Transformer-based architectures, can exceed 400 million parameters. Typically, they are first pre-trained on large-scale EEG data from diverse BCI tasks and later fine-tuned for specific downstream applications. These models demonstrate competitive performance across various downstream tasks while maintaining a unified architectural framework (Jiang et al. 2024; Yue et al. 2024; Wang et al. 2025). While effective, this pre-training paradigm is computationally expensive, often requiring significant GPU time, storage, and energy (Jiang et al. 2024). The high cost also limits the feasibility of further exploration into parameter optimization and neural architecture search (Yu et al. 2023).
@@
Recent advances in computer vision have shown that it is possible to significantly reduce the computational burden of large-scale model training by distilling extensive datasets into compact, information-rich subsets (Zhang et al. 2024; Joshi et al. 2024; Killamsetty et al. 2021). Motivated by the broader goal of data-efficient pre-training, we turn our focus to EEG data, which presents unique and domain-specific challenges that demand tailored solutions. First, EEG signals exhibit a low signal-to-noise ratio (SNR), as neural activity is easily obscured by artifacts from eye movements, muscle activity, and external interference (Sadiya et al. 2021). Even after preprocessing, residual noise can persist and adversely impact model learning. Second, EEG recordings often contain significant redundancy; temporally adjacent segments may capture overlapping or repetitive patterns, offering limited new information (Amin et al. 2016). These characteristics make EEG fundamentally different from vision data in structure and signal quality, necessitating data-efficient strategies that directly address its inherent noise and redundancy. Despite these challenges, little is known about how the volume and composition of pre-training EEG data influence the generalization ability of foundation models, leaving a critical gap in the development of scalable and effective large-scale EEG modeling.
To overcome these limitations, we propose EEG-DLite, a data distillation framework for data-efficient pre-training of large EEG foundation models. EEG-DLite systematically prunes large, unlabeled EEG datasets that are collected using varied channel montages into smaller, more representative subsets. Due to the high dimensionality and low signal-to-noise ratio of EEG signals, instead of selecting samples directly using EEG signals, our approach begins by encoding EEG segments into compact latent representations using a self-supervised autoencoder. Based on these representations, we apply two key techniques:
@@
These steps produce a significantly smaller dataset that still captures the diversity and structure necessary for effective model training. In our experiments, training on only 5% of a 2,500-hour EEG dataset distilled using EEG-DLite achieves comparable, or even superior, performance on several downstream tasks compared to training on the full dataset, as shown in Figure 1. Additionally, this reduces GPU pre-training time from 30 hours to just 2 hours, under the same hardware conditions.
To our knowledge, this is the first study to explore data distillation for physiological signals like EEG in the context of large foundation models. We also perform systematic comparisons of generative and selection-based distillation approaches, providing new insights into what makes pre-training data effective in EEG modeling [^1] The code is available at https://github.com/t170815518/EEG-DLite.
@@
  • We present the first comparison between synthetic data generation and selection-based distillation on EEG. The results show the challenges EEG poses on the generation approach and highlight the effectiveness of a core-set selection approach.
@@
4: Compute outlier scores: $\sum_{i=1}^{d}\log\frac{1}{p_{i}(x_{i})+\alpha}$
@@
[…]
7: Return subset $\mathcal{S}\subset\mathcal{X^{\prime}}$
@@
Figure 2: Overview of the proposed framework. In the first stage, an auto-encoder is trained with self-supervised learning to project each EEG segment $X$ into a lower-dimensional latent representation $\mathbf{z}$. In the second stage, out-of-distribution scores are computed for all samples, and the top $\tau\%$ with the highest scores are excluded. Finally, $\eta\%$ of the datasets is selected via diversity sampling to retain only the most informative data.
EEG foundation models
@@
EEG foundation models have been an emerging paradigm for decoding and modeling brain activity using deep neural networks pretrained on large-scale EEG recordings (Zhou et al. 2025; Lai et al. 2025). Typical architectures adopt self-supervised learning (SSL) objectives and employ CNNs, transformers, or hybrid architectures to extract generalizable neural representations. These models have shown strong performance on downstream tasks such as cognitive state decoding and clinical classification (Jiang et al. 2024; Yue et al. 2024). However, training them requires massive datasets, often exceeding millions of EEG segments, which poses significant computational and storage burdens. More critically, existing efforts largely overlook the influence of training data composition and selection on model performance. While prior work has focused on architectural innovation and transfer learning strategies, few have explored how data quality, representativeness, and redundancy affect the foundation model’s generalization ability. Our work addresses this gap by introducing a data-centric framework that selects diverse, informative EEG segments to reduce training costs without compromising generalization. By emphasizing principled data selection, we aim to enhance both the scalability and explainability of large EEG models, bridging the methodological gap between neuroscience and AI through efficient and standardized data curation.
Data distillation
@@
Data distillation refers to the process of constructing a compact and informative subset of training data that allows a model to achieve comparable performance to that trained with the full dataset, while significantly reducing computational costs. Existing distillation methods generally fall into two main categories: synthetic data generation and core-set selection (Joshi et al. 2024; Sener and Savarese 2018). Synthetic data generation focuses on generating samples that simulate the learning behavior or statistical properties of the original dataset. They have gained popularity in many computer vision tasks due to the flexibility and expressive capacity (Ding et al. 2024). These methods often use optimization objectives such as performance matching, parameter matching, or distribution matching (Yu et al. 2023; Zhang et al. 2024). However, they typically rely on bi-level optimization, which involves nested optimization loops, making them computationally intensive. Core-set selection, in contrast, selects a real and representative subset from the training data (Sener and Savarese 2018). They are more stable and interpretable since they work with real data, but their selection process becomes increasingly expensive as both data size and dimensionality grow. This is particularly problematic for EEG signals, which are high-dimensional, noisy, and vary across subjects. To address the computational burden, we adopt SSL to train a lightweight autoencoder that compresses EEG segments into low-dimensional latent representations, enabling efficient selection (Balestriero et al. 2023).
@@
This section elaborates on the details of the proposed framework, EEG-DLite, which is a model-agnostic EEG data distillation framework. It is designed to select a compact yet representative subset from massive unlabeled data for efficient pre-training. As illustrated in Figure 2 and Algorithm 1, the framework consists of three main steps: (1) a self-supervised multi-view autoencoder compresses high-dimensional EEG signals into low-dimensional latent representations; (2) an outlier detection module filters out anomalous or low-quality samples that may degrade diversity; and (3) a $k$-center selection algorithm identifies a diverse set of EEG segments that best represent the original dataset.
@@
To handle the high dimensionality of EEG signals, we train a lightweight autoencoder with SSL that maps raw EEG segments into compact latent representations without requiring labeled data (Zhao et al. 2024; Zhang et al. 2023). This architecture is intentionally decoupled from any specific foundation model to ensure flexibility and broad compatibility.
@@
The SSL objective includes two components: a reconstruction loss $\mathcal{L}_{\text{Rec}}$ and an inter-instance discrimination loss $\mathcal{L}_{\text{IDC}}$. The reconstruction loss minimizes the mean squared error between input patches $\bm{x}_{i}$ with length $L$ and reconstructions $\bm{x}_{i}^{\prime}$, ensuring the encoder can accurately encode neural signal content. The loss of each sample $X$ is defined as
@@
$\mathcal{L}_{\text{IDC}}$ penalizes excessive token similarity between samples in one batch, encouraging feature diversity (Chen et al. 2021). Specifically, denoting $\mathbf{z}_{i}$ and $\mathbf{z}_{i}^{\prime}$} the $i$-th token before and after the encoder, the network projects them into the same embedding space with two separate projectors $g_{1}$ and $g_{2}$ and penalizes the cosine similarity between an original token and the encoded tokens in other samples, as described below,
@@
PCA + DS 1 $2.6\pm 0.4 2.2$$\pm 0.5 22$$2.2\pm 0.4$ $356\pm 0.004 0.0$$046\pm 0.005$ $14.90\pm 0.03$
5 $31.0\pm 0.4$ $13.4\pm 0.5$ $31.2\pm 0.4$ $0.534\pm 0.004$ $0.262\pm 0.000$ $13.32\pm 0.06$
10 $36.3\pm 0.3$ $20.1\pm 0.4$ $36.5\pm 0.3$ $\bm{0.541\pm 0.002}$ $0.265\pm 0.003$ $13.26\pm 0.02$
25 $38.6\pm 0.2$ $23.5\pm 0.3$ $39.0\pm 0.2$ $0.546\pm 0.001^{\dagger}$ $0.276\pm 0.004^{\dagger}$ $13.15\pm 0.03^{\dagger}$
Proposed 1 $\bm{29.7\pm 0.3}^{\dagger}$ $\bm{11.3\pm 0.3}^{\dagger}$ $\bm{29.7\pm 0.3}^{\dagger}$ $\bm{0.506\pm 0.007}^{\dagger}$ $\bm{0.225\pm 0.010}^{\dagger}$ $\bm{13.61\pm 0.10}^{\dagger}$
5 $\bm{38.6\pm 0.2}^{\dagger}$ $\bm{23.1\pm 0.2}^{\dagger}$ $\bm{38.9\pm 0.2}^{\dagger}$ $\bm{0.550\pm 0.001}^{\dagger}$ $\bm{0.283\pm 0.001}^{\dagger}$ $\bm{13.15\pm 0.07}^{\dagger}$
10 $\bm{39.1\pm 0.1}^{\dagger}$ $\bm{23.9\pm 0.2}^{\dagger}$ $\bm{39.5\pm 0.2}^{\dagger}$ $\bm{0.541\pm 0.002}$ $\bm{0.277\pm 0.014}$ $\bm{13.07\pm 0.04}^{\dagger}$
25 $\bm{39.7\pm 0.2}^{\dagger}$ $\bm{24.7\pm 0.2}^{\dagger}$ $\bm{40.1\pm 0.2}^{\dagger}$ $\bm{0.550\pm 0.003}^{\dagger}$ $\bm{0.281\pm 0.005}^{\dagger}$ $\bm{13.14\pm 0.06}^{\dagger}$
Full data 100 $41.0\pm 0.6$ $26.1\pm 0.8$ $41.2\pm 0.6$ $0.538\pm 0.010$ $0.288\pm 0.003$ $12.25\pm 0.03$
Table 1: Performance comparison across different distillation ratios ($\eta$) on SEED-V and MoBI datasets. $\bm{\kappa}$ refers to Cohen’s kappa coefficient. Bold values indicate the best performance within each distillation ratio, excluding the full data case. DS refers to the diversity sampling. The symbol indicates the result is significantly better ($p<0.05$) than the corresponding Random baseline at the same $\eta$ based on Mann–Whitney U test.
@@
Each sample $X$ can be assigned with an out-of-distribution (OOD) score, defined as Equation 3. The top $\tau$ percent of samples are excluded from the subsequent diverse selection step to prevent degrading the diversity and quality of the distilled subset.
$\text{OOD}(X)=\sum_{i=1}^{d}\log\frac{1}{p_{i}(x_{i})+\alpha}$ (3)
@@
Table 2: Performance comparison across distillation ratios ($\eta$) on TUEV and TUAB datasets. $\bm{\kappa}$ refers to Cohen’s kappa coefficient. Bold values indicate the best performance within each distillation ratio, excluding the full data case. DS refers to the diversity sampling. The symbol indicates the result is significantly better ($p<0.05$) than the corresponding Random baseline at the same $\eta$ based on Mann–Whitney U test.
The proposed framework utilizes the selection-based distillation method. We formalize the process of selecting the most representative EEG samples as the coreset construction problem. Formally, let $\mathcal{Z}\subset\mathbb{R}^{N\times d}$ be the EEG compressed pre-training dataset, where $N$ is the sample size, $d$ is the dimension of compressed representations. The objective is to select $k$ points from $\mathcal{Z}$ such that every data point $\mathbf{z}$ is closest to one selected center, and the largest distance between any point to its closest center is minimized, which is described as
@@
This problem is NP-hard, so we utilize greedy approximation solver proposed in (Sener and Savarese 2018) to iteratively select points that maximize the minimum distance to previously chosen k-centers, of which the time complexity is $\mathcal{O}(k\times N\times d)$.
Figure 3: UMAP visualization of compressed EEG segments after OOD removal and diversity sampling on SEED-sleep (Li et al. 2025).
@@
This section elaborates on the baselines, implementation details, the setup during pre-training and fine-tuning stages, followed by the results and in-depth analysis.
@@
PCA (Anuragi et al. 2024; Fujiwara et al. 2020) reduces EEG dimensionality by projecting data onto top principal components. We use Incremental PCA to scale efficiently and assess how it compares with SSL-based embeddings in representing EEG diversity.
M3D (Zhang et al. 2024) generates synthetic data by minimizing embedding discrepancies using random, untrained networks. Although it is lightweight and computationally feasible at low distillation ratios, it still becomes impractical at 25 percent distillation ratio due to excessive GPU memory demands.
@@
The input signals from both the potential and spectral domains are independently normalized and segmented into 20 non-overlapping patches. Then, the neural compressor is trained for 50 epochs using the Adam optimizer, with the learning rate of 0.001 and gradient clipping applied at a maximum norm of 5.0. A scheduler is used to reduce the learning rate every 10 epochs with a decay factor of 0.5. The training objective, described in Equation 2, incorporates an instance discrimination loss term $\mathcal{L}_{\text{IDC}}$, weighted by $\beta=0.0001$. The encoder consists of 6 self-attention layers with 8 heads per layer, while the decoder contains 2 transformer layers, each with 8 attention-heads as well. Each EEG segment $X$ is projected into a 64-dimensional embedding space.
@@
We use the same pre-training datasets as in LaBraM (Jiang et al. 2024), which includes 32 datasets over 2,500 hours of EEG data collected from multiple tasks (e.g., emotion recognition, motor imagery, etc.). These datasets have diverse channel montages. Each EEG sample is a segment into $X\in\mathbb{R}^{c\times T}$, where $c$ is the number of channels and $T$ is the temporal length (either 4 or 8 seconds), with a fixed stride of 4 seconds. All data are band-pass filtered within the range [0.1, 75] Hz and resampled to 200Hz. During the distillation process, each dataset is distilled independently.
@@
To assess the effectiveness of each distillation method, we evaluate the pretrained foundation models on diverse types of downstream tasks. To perform rigorous comparison, each method distills the original pre-training dataset at multiple compression ratios ($\eta=1\%,5\%,10\%,25\%$) first, and the resulting distilled subsets are used for pre-training. Finally, the pretrained models are fine-tuned separately on four representative downstream tasks that cover both classification and regression settings. We follow the identical evaluation protocol established in LaBraM to ensure direct comparability. All experiments use consistent hyperparameters during pre-training and fine-tuning. Results are averaged across five random seeds, and standard deviations and results’ significance are reported in Table 1 and 2. Below are the details of downstream tasks:
  • TUEV (Obeid and Picone 2016): Six-class classification of EEG events.
  • TUAB (Obeid and Picone 2016): Binary classification of normal versus abnormal EEG signals.
  • SEED-V (Liu et al. 2022): Five-class emotion recognition based on EEG.
  • MoBI (He et al. 2018): Continuous regression of bilateral lower-limb joint angles from EEG during walking.
@@
The foundation model retains comparable performance with  5% of the pre-training data.
As the dataset percentage increases, the performance on all the downstream tasks demonstrates improvements, indicating enhanced model generalization with more training data, and the performance change is also aligned with the experiments reported in (Jiang et al. 2024). Tables 1 and 2 indicate a consistent pattern of data redundancy in large-scale EEG pre-training datasets. Specifically, we observe that using only  5% of the pre-training data selected by EEG-DLite achieves performance close to the upper bound. In contrast, the random selection baseline requires around 25% of the data to reach a similar level. These results suggest that a substantial portion of the data has limited impact on model performance, highlighting the potential for more efficient training and deployment through targeted data selection.
@@
Method $\bm{\eta}$ (%) Acc. F1 $\bm{\kappa}$
Random 50 53.4 54.0 28.9
25 52.8 52.1 28.0
PCA + DS 50 51.8 48.6 27.4
25 51.6 49.9 26.2
Proposed ($\tau=0$) 50 54.1 52.8 29.9
25 54.6 55.1 29.1
Proposed ($\tau=1\%$) 50 56.6 56.7 33.2
25 55.3 55.7 31.3
Full Data 100 54.6 55.4 30.8
@@
To further evaluate the effectiveness of EEG-DLite, we conduct a pilot study on SEED dataset (Zheng and Lu 2015b; Duan et al. 2013) using a small network like EEGNet in the cross-subject supervised learning setting. The results in Table 3 show that EEG-DLite achieves the highest performance when trained on only 25% of the data, even outperforming models trained on the full dataset. We also observe the model is sensitive to the presence of OOD samples in the supervised setting.